# DurFlex-EVC: Duration-Flexible Emotional Voice Conversion with Parallel Generation

### v3

@@ Abstract
Emotional voice conversion involves modifying the pitch, spectral envelope, and other acoustic characteristics of speech to match a desired emotional state while maintaining the speaker’s identity. Recent advances in EVC involve simultaneously modeling pitch and duration by exploiting the potential of sequence-to-sequence models. In this study, we focus on parallel speech generation to increase the reliability and efficiency of conversion. We introduce a duration-flexible EVC (DurFlex-EVC) that integrates a style autoencoder and a unit aligner. The previous variable-duration parallel generation model required text-to-speech alignment. We consider self-supervised model representation and discrete speech units to be the core of our parallel generation. The style autoencoder promotes content style disentanglement by separating the source style of the input features and applying them with the target style. The unit aligner encodes unit-level features by modeling emotional context. Furthermore, we enhance the style of the features with a hierarchical stylize encoder and generate high-quality Mel-spectrograms with a diffusion-based generator. The effectiveness of the approach has been validated through subjective and objective evaluations and has been demonstrated to be superior to baseline models.
@@ I Introduction
Emotional voice conversion (EVC) involves modifying various acoustic characteristics of a voice, such as pitch and spectral envelope, to match a desired emotional state while preserving the speaker’s identity. EVC has gained prominence, particularly in the realm of voice-interactive technologies such as virtual assistants and internet of things (IoT) devices, improving the human-like and emotionally resonant aspects of digital interactions [1, 2, 3, 4].
@@
In the context of EVC, a crucial objective is to preserve the speaker identity and content of the original speech while modifying only those speech attributes that convey emotion [5, 6]. This necessitates an adjustment of the prosody to align with the intended emotion. Prosodic elements, including intonation, rhythm, and energy, play a critical role in both conveying and recognizing emotions in speech. Although the concept of controlling prosody for emotional conversion is intuitively appealing [7], the process of refining each prosodic component presents a significant challenge.
@@
The field of EVC has been revolutionized by advances in deep learning [8, 9]. Some studies employed Gaussian mixture models [10] to convert spectral and prosodic features to produce more expressive voices. Subsequent developments led to autoencoder-based methods [11, 12, 13], enabling learning in non-parallel data-driven EVC. To convert non-parallel emotional speech without changing the speaker’s identity and linguistic content, some VAE-based methods [14] have been proposed. GAN-based approaches [15], using frameworks such as Cycle-GAN [16], StarGAN [17], and VAE-GAN [18], represent further advances. However, these methods often overlook the importance of rhythm when expressing emotion because they support emotional conversion with a fixed length.
@@
Sequence-to-sequence (Seq2Seq)-based models, capable of implicitly modeling duration, have emerged as a significant development [19, 20]. These models often adopt specific strategies, such as a two-stage learning strategy integrating a text-to-speech (TTS) model, to improve learning stability [21, 22]. Although seq2seq models can generate varying durations, they face typical autoregressive model challenges, such as long-term dependency and repetition issues. This necessitates a parallel generation approach for efficiency and reliability. In this study, we explored parallel generation methods with flexible durations. To enable parallel generation, the duration of the content needs to be explicitly modeled. Some voice conversion models [23] have leveraged phoneme duration from text-to-speech models. Obtaining phoneme durations requires additional effort, such as using the encoder-decoder attention of a pre-trained autoregressive TTS or using an external forced alignment tool.
Recently, the exploration of discrete speech units through self-supervised learning representations has shown promise in addressing parallel generation challenges in speech processing. This technique encodes speech into discrete units, facilitating frame-level duration extraction for each unit. Certain studies [24, 25] have investigated leveraging these properties. For instance, [24, 25] proposed an approach similar to spoken language translation for speech emotion conversion, allowing parallel audio generation by predicting the duration of each unit. However, this approach does not fully achieve parallel generation due to its reliance on autoregressive models for emotion translation.
In this paper, we propose a duration-flexible EVC (DurFlex-EVC) that supports parallel generation. We consider discrete speech units as the key to parallel generation and duration flexibility. Although the units have content information, the acoustic characteristics are not sufficient, and thus we use other speech representations as input instead. Flexible duration is achieved by guiding the model to predict units, deduplicating features into unique units, and scaling to frame level by predicting unit durations corresponding to target emotions. Our main contributions are as follows.
  • We achieved flexible duration modeling using discrete speech units and introduced unit aligners to model emotional stylistic context.
  • We designed a style autoencoder to disentangle the content and emotional style of input features. The style autoencoder de-stylizes source emotion from input features and stylizes target emotion for the desired emotion.
  • We introduced a hierarchical stylize encoder that adapts styles at the unit-level and frame-level.
  • We adopted a diffusion-based generator to produce high-quality speech.
  • We propose emotion embedding cosine similarity (EECS) as a method to objectively evaluate the emotional expressiveness of generated speech.
  • We conduct comprehensive evaluations to demonstrate the effectiveness of each component and their contributions to the overall performance of the model.
@@ II-A Exploring Self-Supervised Learning in Speech
Self-supervised learning (SSL) is a machine learning paradigm in which models are trained on their own datasets to create meaningful representations. This method is particularly beneficial in speech processing, where data labeling can be both time-consuming and costly. The wav2vec 2.0 model [26] used contrastive learning to validate SSL representations. The vq-wav2vec model [27] introduced a technique to learn discrete audio representations through self-supervised context prediction and quantization. XLS-R [28] is an extensive cross-lingual speech representation model based on wav2vec 2.0. Hidden-unit BERT (HuBERT) [29] employs a masked prediction approach similar to BERT [30] for learning representations. ContentVec [31] improves speaker disentanglement within the HuBERT framework. Efforts have also been made to create representations suitable for universal downstream tasks [32].
Recently, SSL representations have been extensively applied to various downstream tasks, such as automatic speech recognition [26], voice conversion [33], speaker verification [34], speech synthesis [35], speech emotion recognition [36], and speech enhancement [37].
@@ II-B Discrete Units in Speech Processing
In the realm of audio and speech, discrete unit representation has been proposed for diverse tasks. SoundStream [38] introduced a neural audio codec employing a residual vector quantizer (RVQ), while EnCodec [39] focuses on high-fidelity audio compression and lightweighting through similar methods. UniAudio [40] emerged as a general-purpose audio generation model. These neural codec-based methods, aimed primarily at audio compression and restoration, feature large codebooks and relatively small dimensions.
In contrast, certain methods emphasize the compression of speech into semantic units. A method was proposed to decompose and reconstruct speech into discrete units of pitch and speaker identity [41]. Based on this, soft speech units were suggested [24] for enhanced content capture, thereby improving the naturalness and intelligibility. Furthermore, speech emotion conversion was explored as a language translation task [42], using discrete representations of phonetic content, prosody, speaker, and emotion in conjunction with a neural vocoder for waveform generation. UnitSpeech [25] demonstrated proficiency in personalized TTS and voice conversion, fine-tuning a diffusion-based TTS model with minimal data using self-supervised units, eliminating the need for retraining for each task. Here, semantic speech units serve as content in speech decomposition.
@@ II-C Parallel Speech Generation
Non-autoregressive speech synthesis models generate speech frames in parallel, significantly reducing inference time compared to autoregressive methods. In a recent speech synthesis study, parallel generation methods outperformed and were more reliable than autoregressive methods for text-to-speech, voice conversion, and vocoder. For parallel generation, TTS requires an alignment between text and speech, which can be extracted from pre-trained autoregressive teacher models [43], using an external aligner [44] such as the Montreal Forced Aligner (MFA) [45], or using the monotonic alignment search algorithm (MAS) [46]. For voice conversion, many approaches [47, 48] employ parallel generation, but some models adopt autoregressive [49] due to the impossibility of converting duration. Some works [50, 23] have proposed models with variable duration and parallel generation; these models use text-speech alignment as in TTS. For vocoders, studies have been proposed to upsample the input features and then generate waveforms through various generation models such as GAN [51], normalizing flow [52], or generative diffusion model[53]. Inspired by the success of parallel generation models, we aimed to adopt a parallel generation framework for emotional voice conversion.
@@ II-D Duration Modeling in Speech Processing
Duration modeling is a critical aspect of speech synthesis, particularly in TTS, where mismatches between character length and signals can occur. Early TTS methods addressed this through autoregressive models with implicit duration modeling, generating one frame at a time [54, 55]. FastSpeech [43] leveraged the encoder-decoder attention alignment of an autoregressive teacher model to model phoneme duration, facilitating parallel generation. FastSpeech 2 [44] introduced a method for extracting phoneme duration from forced external alignment. Glow-TTS [46] developed a technique for identifying the most similar monotonic alignment between text and latent using a monotonic alignment search. Duration modeling advances have also been applied to voice conversion, with seq2seq models handling duration changes [56, 49]. The DCVC model [23] utilized a phoneme-based information bottleneck for style transfer and speech speed control in voice conversion. Discrete speech units have been employed [41, 25] to model duration through consecutive unit counts. In EVC, a trend towards seq2seq structures for handling duration changes has been observed [22, 57]. However, despite [42] proposing EVC using units for parallel generation, it still relies on the seq2seq model for the unit translation process. We leveraged units to model duration without alignment to text-speech to facilitate a variable-length parallel generation framework.
Figure 1: Overall framework of the proposed method. The feature extractor transforms the source audio into input features. These features are subsequently disentangled and reconditioned by the style autoencoder. The unit aligner is responsible for providing unit-level context and performing duration modeling. In addition, the hierarchical style encoder encodes features at both the unit and frame levels. Mel-spectrograms are subsequently produced by the generator. In this figure, ”DP” represents the duration predictor, $\mathcal{LR}$ denotes the length regulator, while ”Q”, ”K” and ”V” represent the query, key and value of the cross-attention in the unit aligner, respectively. $\otimes$ denotes the concatenate operation. $w_{src}$ represents the source style vector and $w_{tgt}$ represents the target style vector. The style autoencoder disentangles the source style from the features and applies the target style, while the hierarchical stylize encoder and generator take the target style as a condition.
@@ III Proposed Method
We propose DurFlex-EVC, which is flexible in duration and parallel in generation. The model consists of the following components:
  • Feature extractor that transforms raw audio waveforms into acoustic features.
  • Style autoencoder, facilitating the disentanglement of content and style.
  • Unit aligner that handles context transformation and precise duration modeling.
  • Hierarchical stylize encoder that operates at both unit and frame levels to capture stylization.
  • Diffusion-based generator that produces high-quality Mel-spectrograms.
Fig. 1 illustrates the comprehensive framework of DurFlex-EVC, with detailed explanations provided in the following subsections.
@@ III-A Overview
Fig. 1a shows the overall structure of the model. First, the waveform is transformed by the feature extractor to be used as input to the model. Features such as Mel-spectrograms or SSL representations can be used as input. In our method, we adopted HuBERT [29] as a feature extractor and used the output of the last layer as input to the model. The style of the input feature is disentangled and is adapted by the style autoencoder. The unit aligner then aggregates contextual information at the frame level through a cross-attention module. This representation is compressed to the unit level, feeding into the hierarchical stylize encoder. Features are stylized at the unit-level, and then stylized after frame-level extension. The diffusion-based generator produces a Mel-spectrogram from the output of the hierarchical stylize encoder and a style vector. This Mel-spectrogram is then converted into a raw waveform by a pre-trained vocoder.
@@ III-B Style Autoencoder
The feature extractor generates features that include both content and style. In this context, content refers to the linguistic information of the speech, while style encompasses all other aspects, including emotional expression and speaker-specific properties. To model styles, we distinguish between speaker style and emotional style. The style autoencoder is designed to decompose the source emotion style. Fig. 1b shows the structure of the style autoencoder, which consists of two primary components: the de-stylize transformer and the stylize transformer. The de-stylize transformer causes the style to be decomposed from the feature, while the stylize transformer applies the style. The de-stylize transformer is shown in Fig. 1f and the stylize transformer in Fig. 1g. These components have been designed using advanced normalization techniques.
Firstly, layer normalization (LN) serves as a fundamental technique, mathematically expressed as follows:
@@
where $z$ represents the input vector to be normalized, $\mu$ is the mean of the vector, and $\sigma$ is its standard deviation.
The stylize transformer employs conditional layer normalization (CLN) [58] based on LN to effectively adapt styles. CLN is defined as follows:
@@
where $\gamma(w)$ and $\beta(w)$ are adaptive parameters representing the gain and bias for the style vector $w$.
In contrast, the de-stylize transformer employs mix-style layer normalization (MixLN) [59], an modification of CLN, to disentanlge style-independent features. MixLN introduces perturbations in the input directed towards the style vector, inhibiting the model’s tendency to learn style-specific features. This perturbation is executed by blending the original style vector with batch-level shuffled style vector.
$\gamma_{mix}(w)=\lambda\gamma(w)+(1-\lambda)\gamma(\tilde{w}),$ (3)
$\beta_{mix}(w)=\lambda\beta(w)+(1-\lambda)\beta(\tilde{w}),$ (4)
where $w$ and $\tilde{w}$ denote the original and shuffled style vector, respectively. The $\lambda$ is responsible for balancing the original style with the shuffle style and follows a beta distribution, $\text{Beta}(\alpha,\alpha)$ with $\alpha\in(0,\infty)$. This parameter resides in a $B$-dimensional real number space, denoted as $\mathbb{R}^{B}$, where $B$ represents the batch size, the amount of data processed simultaneously by the model. Therefore, MixLN is defined as follows:
$\text{MixLN}(z,w)=\gamma_{mix}(w)\times\text{LN}(z)+\beta_{mix}(w).$ (5)
A style autoencoder is composed of $N$ de-stylize transformers and $N$ stylize transformers. We constructed the source style vector $w_{src}$ as the sum of the speaker vector $s_{src}$ and the emotion vector $e_{src}$, and the target style vector $w_{tgt}$ as the sum of the speaker vector $s_{src}$ and the emotion vector $e_{tgt}$.
@@
The speaker vector $s_{*}$ and the emotion vector $e_{*}$ are obtained from the embedding look-up table. The de-stylize transformer uses $w_{src}$ to disentangle the source emotion style from input feature, whereas the stylize transformer utilizes $w_{tgt}$ to apply the target emotion style.

### v4

# DurFlex-EVC: Duration-Flexible Emotional Voice Conversion Leveraging Discrete Representations without Text Alignment

@@ Abstract
Emotional voice conversion (EVC) involves modifying various acoustic characteristics, such as pitch and spectral envelope, to match a desired emotional state while preserving the speaker’s identity. Existing EVC methods often rely on text transcriptions or time-alignment information and struggle to handle varying speech durations effectively. In this paper, we propose DurFlex-EVC, a duration-flexible EVC framework that operates without the need for text or alignment information. We introduce a unit aligner that models contextual information by aligning speech with discrete units representing content, eliminating the need for text or speech-text alignment. Additionally, we design a style autoencoder that effectively disentangles content and emotional style, allowing precise manipulation of the emotional characteristics of the speech. We further enhance emotional expressiveness through a hierarchical stylize encoder that applies the target emotional style at multiple hierarchical levels, refining the stylization process to improve the naturalness and expressiveness of the converted speech. Experimental results from subjective and objective evaluations demonstrate that our approach outperforms baseline models, effectively handling duration variability and enhancing emotional expressiveness in the converted speech.
@@ I Introduction
Emotional voice conversion (EVC) involves modifying various acoustic characteristics of a voice, such as pitch and spectral envelope, to match a desired emotional state while preserving the speaker’s identity[1]. EVC has gained prominence, particularly in the realm of voice-interactive technologies such as virtual assistants and internet of things (IoT) devices, improving the human-like and emotionally resonant aspects of digital interactions [2, 3, 4, 5].
@@
In the context of EVC, a crucial objective is to preserve the speaker identity and content of the original speech while modifying only those speech attributes that convey emotion [1, 6]. This necessitates an adjustment of prosody to align with the intended emotion. Prosody elements, including intonation, rhythm, and energy, play a critical role in both conveying and recognizing emotions in speech. Although the concept of controlling prosody for emotional conversion is intuitively appealing [7], refining each prosody component presents a significant challenge.
@@
The field of EVC has been revolutionized by advances in deep learning [8, 9]. Initial approaches employed Gaussian mixture models [10] to convert spectral and prosody features, producing more expressive voices. Subsequent developments led to autoencoder-based methods [11, 12, 13], enabling learning in non-parallel data-driven EVC. VAE-based approaches [14] and GAN-based frameworks [15]—such as Cycle-GAN [16], StarGAN [17], and VAE-GAN [18]—represent further advances. However, these methods often overlook the importance of rhythm in expressing emotion, as they typically support emotional conversion with fixed durations.
@@
Sequence-to-sequence (Seq2Seq) models, capable of implicitly modeling duration, have become a notable development [19, 20]. These models often employ strategies, such as a two-stage learning approach integrating a text-to-speech (TTS) model, to improve stability [21, 22]. However, Seq2Seq models face challenges like long-term dependency and repetition, necessitating parallel generation for efficiency and reliability. Parallel generation requires explicit duration modeling, which many voice conversion models [23] achieve through phoneme duration derived from TTS alignments. Obtaining phoneme duration often requires encoder-decoder attention from pre-trained autoregressive TTS models[24] or external forced alignment tools[25].
Recently, discrete speech units derived from self-supervised learning representations have shown promise in addressing parallel generation challenges [26, 27, 28, 29]. Certain studies [28, 29] have proposed methods that model speech emotion conversion as a translation task, thereby enabling parallel audio generation. However, these methods still rely on autoregressive models for emotional translation, limiting their effectiveness in fully parallel generation.
In this paper, we propose a duration-flexible EVC framework, DurFlex-EVC, that eliminates the need for text or alignment information while supporting parallel generation. We use discrete speech units to model content and incorporate self-supervised representations to enrich acoustic details. Our model predicts unit sequences and their durations, which allows for flexible duration control.
Our main contributions are as follows:
  • We propose DurFlex-EVC, a duration-flexible EVC framework that uses discrete speech units for content modeling, eliminating the need for text or external alignment information while supporting efficient parallel generation.
  • We introduce a unit aligner that models contextual relationships between speech features and unit sequences, enabling effective duration control without relying on text-based alignment.
  • We develop a style autoencoder that separates and reintroduces emotional styles in the input features, facilitating content-style disentanglement for emotional voice conversion.
  • We design a hierarchical stylize encoder that captures both global and local emotional patterns, thereby enhancing the expressiveness of emotional speech.
  • We perform extensive subjective and objective evaluations, demonstrating the superior performance of our method compared to existing approaches.
@@ II-A Exploring Self-Supervised Learning in Speech
Self-supervised learning (SSL) is a machine learning paradigm in which models are trained on unlabeled datasets to generate meaningful representations. This approach is particularly advantageous in speech processing, where labeling data is labor-intensive and expensive. Models like wav2vec 2.0 [30] and vq-wav2vec [31] have demonstrated SSL’s potential, with wav2vec 2.0 leveraging contrastive learning and vq-wav2vec introducing quantization techniques for discrete representations. Subsequent developments, such as XLS-R [32], extend these ideas to cross-lingual applications, while models like Hidden-unit BERT (HuBERT) [33] and ContentVec [34] focus on masked prediction and speaker disentanglement, respectively.
SSL representations have proven valuable across a wide range of speech-related tasks, including automatic speech recognition [30], voice conversion [35], speaker verification [36], and speech emotion recognition [37]. The increasing adoption of SSL techniques motivates further exploration into how these representations can be leveraged for more complex and resource-efficient speech applications, such as emotional voice conversion, where conventional supervised methods are often limited by data availability.
@@ II-B Discrete Units in Speech Processing
Discrete unit representations have emerged as a versatile tool in audio and speech processing [38, 39, 40]. Techniques like SoundStream [38] and EnCodec [39] focus on high-fidelity audio compression, employing methods such as residual vector quantization (RVQ), while UniAudio [40] serves as a general-purpose audio generation model. These neural codec-based methods, aimed at audio compression and restoration, utilize large codebooks and compact dimensions. Although these approaches excel in compression, they are less suited for tasks that require detailed modeling of linguistic and prosodic content.
In contrast, some methods emphasize encoding speech into semantic units to facilitate applications like speech synthesis [29] and emotion conversion [33]. For instance, techniques have been developed to decompose and reconstruct speech into discrete units for content, pitch, and speaker identity [26]. To improve naturalness and intelligibility, soft speech units were introduced [28] for enhanced content capture. Speech emotion conversion has also been framed as a language translation task [27], leveraging discrete representations of phonetic content, prosody, speaker, and emotion alongside neural vocoders for waveform generation. UnitSpeech [29] has demonstrated effectiveness in personalized TTS and voice conversion, using self-supervised units to fine-tune a diffusion-based TTS model with minimal data, thus bypassing the need for retraining across tasks. This focus on discrete units offers a promising solution to the challenges of aligning speech content, offering the potential for more efficient and accurate modeling.
@@ II-C Parallel Speech Generation
Non-autoregressive speech synthesis models generate speech frames in parallel, significantly reducing inference time compared to autoregressive methods [24, 41, 42]. Recent studies have demonstrated that parallel generation methods can offer superior performance and reliability over autoregressive approaches for applications like text-to-speech, voice conversion, and vocoders. In text-to-speech systems, parallel generation typically relies on precise alignments between text and speech. These alignments are often derived using pre-trained autoregressive teacher models [24], external aligners like the Montreal Forced Aligner (MFA) [43], or algorithms such as the monotonic alignment search (MAS) [41]. Voice conversion methods often attempt to leverage the benefits of parallel generation. While many approaches [44, 45] have adopted parallel generation strategies, certain models remain autoregressive [46] due to complexities in handling variable durations. To address these challenges, some solutions [47, 23] have integrated variable-duration modeling with parallel generation. However, this often requires text-speech alignment similar to TTS frameworks. Additionally, vocoders have used techniques like GANs [48], normalizing flows [49], and generative diffusion models [50] to efficiently produce high-quality waveforms. The success of parallel generation frameworks has inspired further exploration into adapting these techniques for emotional voice conversion, with the goal of achieving efficient and expressive speech synthesis without relying on traditional alignment methods.
@@ II-D Duration Modeling in Speech Processing
Duration modeling is crucial in speech synthesis, particularly in TTS systems, where discrepancies between character length and audio signals often occur [24]. Early approaches used autoregressive models to implicitly handle duration, generating speech one frame at a time [51, 52]. FastSpeech [24] improved upon these methods by leveraging alignments from an autoregressive teacher model to explicitly model phoneme duration, thereby enabling parallel generation. Building on this, FastSpeech 2 [25] introduced phoneme duration extraction using external forced alignment. Glow-TTS [41] further refined this approach by employing a monotonic alignment search to find optimal matches between text and latent representations. Duration modeling has also been explored in voice conversion, where seq2seq models are often used to manage variable durations [53, 46]. The DCVC model [23] utilized a phoneme-based information bottleneck for style transfer and speech speed control in voice conversion. Discrete speech units have also been leveraged in voice conversion for modeling duration by counting consecutive occurrences of each unit. This approach allows for effective duration control while bypassing the need for explicit text or phoneme alignment [26, 29]. In EVC, seq2seq structures have become a popular choice for managing duration variations [22, 54]. [27] leveraged discrete units for parallel emotional voice conversion but still incorporated seq2seq models for the translation of unit sequences, introducing challenges inherent to autoregressive architectures. Our work aims to address these limitations by employing discrete units to model duration, eliminating the need for text-speech alignment and enabling a more flexible parallel generation framework.
Figure 1: Overall framework of the proposed method. The feature extractor transforms the source audio into input features. These features are subsequently disentangled and reconditioned by the style autoencoder. The unit aligner is responsible for providing unit-level context and performing duration modeling. In addition, the hierarchical style encoder encodes features at both the unit and frame levels. Mel-spectrogram is subsequently produced by the generator. In this figure, ”DP” represents the duration predictor, $\mathcal{LR}$ denotes the length regulator, while ”Q”, ”K” and ”V” represent the query, key and value of the cross-attention in the unit aligner, respectively. $\otimes$ denotes the concatenate operation. $w_{src}$ represents the source style vector and $w_{tgt}$ represents the target style vector. The style autoencoder disentangles the source style from the features and applies the target style, while the hierarchical stylize encoder and generator take the target style as a condition.
@@ III Proposed Method
Our work introduces DurFlex-EVC, a duration-flexible and parallel generation framework for emotional voice conversion. We designed this system to overcome limitations of previous approaches, such as the reliance on seq2seq structures and the need for external text or alignment information. The key components of our model and their roles are as follows:
  • Feature extractor: Transforms raw audio waveforms into acoustic features.
  • Style autoencoder: Designed to separate the emotional style from input features and apply the target emotional style, enabling content-style disentanglement without seq2seq dependencies.
  • Unit aligner: Converts stylized features into unit-level content, capturing contextual information and handling duration modeling, which allows for efficient parallel context transformation.
  • Hierarchical stylize encoder: Applies emotional styles at both unit and frame levels, ensuring that stylistic nuances are captured effectively.
  • Diffusion-based generator: Produces high-quality Mel-spectrogram, which is then converted into speech by a pre-trained vocoder.
@@ III-A Overview
Fig. 1a illustrates the overall architecture of DurFlex-EVC. The process begins with the feature extractor. It transforms raw audio waveforms into acoustic features. While various feature representations, such as Mel-spectrogram or SSL outputs, can be utilized, our approach employs the final layer representations from HuBERT [33]. This choice provides continuous representations that retain comprehensive acoustic information, essential for accurately modeling emotional characteristics in speech. Next, the style autoencoder separates the emotional style from the input features and applies the desired target emotion. This disentanglement ensures that content and style are managed independently, allowing for precise emotional adaptation without relying on seq2seq structures. The unit aligner then processes the stylized features by aggregating contextual information at the frame level through a cross-attention module. This aggregated representation is compressed to the unit level, preparing it for further processing by the hierarchical stylize encoder. The hierarchical stylize encoder operates at both unit and frame levels, refining the emotional style to ensure consistency and expressiveness across different scales. Finally, the diffusion-based generator synthesizes a high-quality Mel-spectrogram from the output of the hierarchical stylize encoder and the style vector. This Mel-spectrogram is subsequently converted into a raw waveform using a pre-trained vocoder, completing the emotional voice conversion process.
@@ III-B Style Autoencoder
The feature extractor generates representations that encompass both content and style aspects of the speech. In this context, content refers to the linguistic information, while style includes emotional expression and speaker-specific characteristics. To model styles, we distinguish between speaker style and emotional style. The style autoencoder is designed to separate the source emotional style from the input features and apply a target emotional style, facilitating content-style disentanglement. Fig. 1b shows the structure of the style autoencoder, which consists of two primary components: the de-stylize transformer and the stylize transformer.
Layer Normalization (LN) is employed as a fundamental technique in both transformers, defined as:
@@
where $z$ represents the input vector, $\mu$ is the mean, and $\sigma$ is the standard deviation. LN ensures that the inputs to each layer have a consistent distribution, facilitating stable and efficient training.
De-stylize Transformer: This component removes the source emotional style from the input features, isolating the content. It employs mix-style layer normalization (MixLN) [55], which blends the original style vector with a shuffled style vector to disrupt the association between features and the source style, as shown in Fig. 1f. The MixLN is defined as:
$\text{MixLN}(z,w)=\gamma_{mix}(w)\times\text{LN}(z)+\beta_{mix}(w),$ (2)
where $\gamma_{mix}(w)=\lambda\gamma(w)+(1-\lambda)\gamma(\tilde{w})$ and $\beta_{mix}(w)=\lambda\beta(w)+(1-\lambda)\beta(\tilde{w})$. $w$ is the original style vector, $\tilde{w}$ is the shuffled style vector, and $\lambda$ is a parameter sampled from a Beta distribution, $\text{Beta}(\alpha,\alpha)$. This process helps in disentangling style-independent content by preventing the model from encoding style-specific features.
Stylize Transformer: This component applies the target emotional style to the content features. It employs conditional layer normalization (CLN) [56], which adjusts the normalized input based on the target style vector $w$, as illustrated in Fig. 1g. The CLN is defined as:
@@
where $\gamma(w)$ and $\beta(w)$ are adaptive parameters derived from the style vector $w$. CLN allows the model to incorporate the desired emotional style into the content features seamlessly.
The style autoencoder consists of $N$ de-stylize transformers and $N$ stylize transformers. We construct the source style vector $w_{src}$ by combining the speaker vector $s_{src}$ and the emotion vector $e_{src}$, and the target style vector $w_{tgt}$ by combining $s_{src}$ with the desired emotion vector $e_{tgt}$:
@@
Both the speaker vector $s_{*}$ and the emotion vector $e_{*}$ are obtained from the embedding look-up table. The de-stylize transformer uses $w_{src}$ to disentangle the source emotional style from the input features, while the stylize transformer applies $w_{tgt}$ to encode the target emotional style into the features. This design enables the model to adapt emotional styles effectively without relying on seq2seq structures, thereby supporting parallel processing and flexible duration modeling.